On his 40th birthday ofentse decides he will buy a bike trailer for his 50th birthday. he estimates that it will cost him 48 000 when he turns 50.he starts saving immediately each month paying an amount into an account earning 8.58% interest per year ,compounded monthly the monthly amount payment is


1.592,95

2.597,19

3. 252,18

4. 253,99


1
Expert's answer
2022-03-20T06:41:44-0400

A=p(1+r/n)nt

48000=p(1+8.58%/12)10*12

48000=p(1.00715)120

48000=2.35125p

P=20414.67

Ofentse's yearly payment is 20414.67

So the monthly payment is 20414.67/12

=1701.22